Find derivative using 1st principle:
f (x)=x1/2-1/x½
 
 
 

Grade:

1 Answers

unnati kasaundhan
6 years ago
f`(x)= lim f (x+h) - f (x) h->0 ------------- h It is the formula of first principle of differential calculus from which you can solve your question by putting following values in the upper formula. .......1) f (x) = x^1/2 -1/x^1/22) f (x+h)= (x+h)^1/2 - 1/(x+h)^1/23) then solve it 4) now follow limit rules to find your answer. ......That`s the way to solve differential problem using first principle ..I hope you can solve your question. ...it is enough to understand. ..So approve my answer...........Ok by dood andkeep practicing